Technical Staff Blog

Last update on .

Max diagnosed and replaced a broken HDMI cable running between the amplifier and the projector in 316, so the AV system should be back at full strength. 
If you experience issues with this room, or AV in any of rooms, please make sure and email so we can work to address ...

Last update on .

CIS will be performing maintenance on their DNS and DHCP services on February 10th between 6am and 7am. They anticipate a degradation of services during this window, which means you may experience issues:

  • reaching Brown services, outside the department
  • obtaining a DHCP address on the RLAB networks, wired and wireless

Yearly archives